# Efficient Bayesian analysis of kilonovae and GRB afterglows with fiesta

### v1

@@ 3.3 KN surrogates
Figure 3: Benchmarks of the surrogates for the KN possis model. We show the deviations of surrogate predictions against a test data set of size $n_{\text{test}}=2238$. Figure layout as in Fig. 1. The figure compares two different surrogates, one using the MLP architecture (blue), the other a is a LightcurveModel, where an MLP is trained for each bandpass separately (green).
@@
$Y_{e,\text{dyn}}=a\cos(\theta)^{2}+b\ ,$ (7)
where $a=0.71b$ (Setzer et al., 2023) and b is scaled to achieve the desired mass-averaged electron fraction $\bar{Y}_{e,\text{dyn}}$. The electron fraction for the wind ejecta, however, is assumed to be constant and thus we do not mark its symbol $Y_{e,\text{wind}}$ with a bar.
We trained three possis surrogates, two as a FluxModel with MLP and cVAE architecture respectively, the other as a LightcurveModel. These surrogates are based constitute an update for the Bu2019 surrogate implemented in nmma that was based on an older possis version. Our training data consists of $n_{\text{train}}=17899$ flux densities that were calculated from $10^{7}$ photon packages. Parameters for the training data set are randomly drawn in the ranges from Table 1, though we note that the ejecta masses were drawn uniformly, not in log-space but in linear space. Furthermore, the inclination $\iota$ is confined to a fixed grid from the possis output that is spaced uniformly in $\cos(\iota)$ between 0 (edge-on) and 1 (face-on). We also mention that the LightcurveModel additionally receives the redshift $z$ sampled randomly between 0 and 0.5 as a training parameter, so that the surrogate automatically incorporates the K-correction of the filter magnitudes. Finding no significant performance difference for various setups, we set the number of PCA coefficients for the MLP architecture in the FluxModel and LightcurveModel to 100, the cVAE was trained on a down-sampled flux array of dimensions $64\times 40$. Training took about 4 minutes for the MLP architectures and 27 minutes for the cVAE on an H100 GPU.
In Fig. 3 we benchmark the surrogate models for the possis KN model. In particular, we compare the MLP FluxModel against the MLP LightcurveModel, since the FluxModel with the cVAE architecture underperforms both. This is because, in contrast to the GRB afterglow training data, the training data from possis contains inherent Monte-Carlo noise, which is more difficult for the cVAE “average out”, as it trains directly on the flux output. The MLP architecture uses the coefficients of the principal components as training input and is thus less receptive to small noise fluctuations. Still, even for those architectures, we generally find higher deviations in the predictions from the test data set than for the afterglow surrogates. In particular, the mismatch rises drastically above 0.5 mag for many test data samples when the flux brightness suddenly drops. This can be seen in Fig. 3, where the mismatch in the right panels is typically confined to $\lesssim 0.5$ mag, but then spikes around the time the KN starts to fade, i.e., after around 1 day in the UV band and after around $\gtrsim$ 10 days in the $i$-band. In general, the dimmer the lightcurve, the higher the Monte-Carlo noise contribution and the larger the mismatch between the surrogate and the test data becomes. However, inspection of random test samples reveals that the surrogate prediction matches quite well at early times, when the absolute magnitude is still brighter than $-10$ mag. At very late times, we truncated the training data below $10^{6.5}$ mJys (corresponding roughly to the 0th absolute magnitude), which causes the mismatch to decrease again when the surrogates pick up on this trend. The LightcurveModel performs slightly better at these later times, however, the FluxModel performs better at earlier times when the emission is still bright. We thus set the MLP FluxModel as the default KN surrogate for the subsequent analyses in Sec. 4 and 5, as the early and bright parts of the lightcurve are most relevant for real-life applications.
We also note that the surrogates from Table 1 are rather extensive in their scope in which they can predict the spectral flux density. Specifically, the GRB afterglow surrogates have been trained across $10^{9}$$5\cdot 10^{19}$ Hz, i.e., from the radio to hard X-ray, and over a time interval of $10^{-4}$$2000$ days. The KN surrogates can predict the expected emission between 0.20.20.20.226262626 days in the frequency range of $10^{14}$$2\cdot 10^{15}$ Hz, i.e., from the far infrared to the far ultraviolet (UV). In certain use cases, the surrogates could easily be finetuned on a smaller frequency or time interval, to potentially deliver even better performance.
4 Bayesian inference with fiesta
Using the surrogate for the determination of the expected light curve $m^{\star}(t,\@vec{\theta})$ in Eq. (2) is an approximation to the real likelihood function. In the present section, we demonstrate that this approximation is still capable of recovering the correct posterior when accounting for the surrogate uncertainty. We do so by using the best-performing surrogates presented in Sec. 3, i.e., the FluxModel instances with cVAE architecture for the afterglow models, and the FluxModel with the MLP architecture for the KN model. We also discuss the performance of the flowMC implementation in fiesta and how it scales when the dimension of the parameter space is increased to include more systematic nuisance parameters in the sampling.
4.1 Injection-recoveries
Figure 4: Parameter recovery of an injected mock lightcurve from the gaussian afterglowpy jet model. The corner plot shows the posterior contours at 68% and 95% credibility. Parameters correspond to the symbols in Table 1, $\sigma_{\text{sys}}$ is the freely sampled systematic uncertainty. Different colors compare posteriors obtained with different sampling methods. The posterior in red is based on likelihood evaluations from the proper afterglowpy model with the nmma sampler. The purple posterior relies on the fiesta surrogate for the likelihood evaluation but uses the nmma sampler. The light blue posterior uses the fiesta surrogate as well but is sampled within fiesta’s own inference framework that relies on flowMC. The injection parameters used to generate the mock lightcurve data are indicated by the orange lines. The insets on the upper right side show the injection data across the photometric filters and the best-fit lightcurve (i.e. highest likelihood) of the fiesta posterior (lightblue) and the nmma posterior (red). The latter lies almost completely underneath the former.
Figure 5: P-P plots for GRB afterglow injections. Each panel shows a P-P plot for the recovery of the parameter displayed in its top left corner. The P-P plots show the cumulative distribution of the injected values’ posterior quantiles for 200 injections. The lightblue curves indicate injection-recoveries with afterglowpy, the magenta ones for pyblastafterglow. The solid lines signify that the injections stem from physical base model, the dashed lines indicate an injection with the surrogate itself. The gray areas mark the 68%-95%-99.7% confidence range in which the quantile distribution should fall if it was uniformly distributed.
To evaluate whether fiesta’s surrogates are capable of recovering the correct posterior, we create mock lightcurve data with the physical base model using randomly drawn model parameters. We then obtain the posterior using the surrogate. The injection data always contain 75 mock magnitude measurements across multiple bands, encompassing the frequency and time range of the model and representing a well-sampled lightcurve. Specifically, for GRB afterglows, the injection data spans from 0.01 to 200 days and contains mock observations from the radio to the X-ray bands. The KN injections reach from the infrared to UV and contain data points between 0.5 to 20 days. For the KN injections, we also apply a detection limit of 24 apparent mag at 40 Mpc, which prevents the surrogates of being used in regions with high prediction error due to high Monte-Carlo noise in the possis training data. We add Gaussian noise to these mock measurements, where the measurement errors $\sigma(t_{j})$ are drawn from a $\chi^{2}$-distribution with one degree of freedom and scaled to lie around $0.1$ mag. To recover the posterior with the surrogate, we use uniform priors across the parameter ranges specified in Tab. 1. The luminosity distance is fixed to $40$ Mpc for each injection.
In Fig. 4 we show how the parameters of one particular injection from the afterglowpy Gaussian jet model are recovered with fiesta, using both PyMultinest in nmma as well as flowMC as samplers, shown in purple and blue, respectively. Since the injection also incorporates a random mock measurement error, the posterior is not always centered around the true injected parameters indicated by the orange lines, but the marginalized posteriors contain the injected values in the 95% credible interval.
Of all the models presented in Sec. 3.2 and 3.3, afterglowpy is the only one for which the execution time of a single likelihood call is sufficiently fast to be used directly in Bayesian inference. Thus, we can compare the approximate posterior obtained with the fiesta surrogate to the posterior obtained using the actual physical base model for the likelihood evaluation. The latter is shown in red in Fig. 4. We find good agreement with the posteriors obtained with the fiesta surrogate, though some small deviations in the posterior tails for the inclination and jet opening angle exist.
We also run four additional afterglowpy injection-recoveries similar to Fig. 4 and compare them to the surrogate posteriors, finding good agreement. For certain degenerate parameters, we even observe that the flowMC sampling algorithm with the surrogate recovers the true injected values better, i.e., the true value lies more at the center of the posterior than with nmma’s PyMultinest sampler using the base model. This can be attributed to broader exploration in the parameter space for degenerate parameters such as $\epsilon_{B}$ and $\log_{10}(n_{\rm ism})$, indicating an advantage in flowMC’s dependence on gradient-based samplers and global proposals.
To systematically evaluate the disagreement between the posteriors caused by the surrogates, we resort to the Kullback-Leibler (KL) divergence $D_{KL}$. Bevins et al. (2025) assessed a theoretical link between the root mean square error (RMSE) of an emulator and the impact on the posterior in terms of $D_{KL}$ if it was obtained with the base model or with the surrogate. For a linear model, they derive the upper-bound on $D_{KL}$
$D_{KL}\leq\frac{N_{d}}{2}\frac{\rm{RMSE}^{2}}{\sigma^{2}}\ ,$ (8)
where $N_{d}$ is the number of data points and $\sigma$ their typical error scale. While this upper-bound is derived by assuming a linear relation $d(\@vec{\theta})$, we can examine this relationship using our afterglowpy injections and the respective surrogate and base model posteriors. We then determine $D_{KL}$ between the posteriors through a kernel density estimate. Using our 5 injection-recoveries with $N_{d}=75$, we find that $D_{KL}$ ranges between 2 – 8 nats and the upper limit is indeed obeyed when setting $\rm{RMSE}=0.1$ mag and taking the largest error from the injected mock data as a conservative value for $\sigma$. In fact, Eq. (8) is a factor of 1.4 – 2.8 above our estimated value for $D_{KL}$. It should be noted however, that this only serves as a limited sanity check, since the bound in Eq. (8) is derived under simplifying assumptions, and determining $D_{KL}$ through kernel density estimates might not be numerically accurate.
We also point out that in order to achieve this agreement between the surrogate posterior and the posterior based on afterglowpy directly, we set a minimum threshold on the systematic uncertainty $\sigma_{\text{sys}}$. Specifically, $\sigma_{\text{sys}}$ was sampled as a free parameter with a uniform prior $\sigma_{\text{sys}}\sim\mathcal{U}(0.3,1)$. Lowering the limit on $\sigma_{\text{sys}}$ can lead to biases in the posteriors recovered with the surrogate, since sampling the systematic uncertainty mainly accounts for potential tension between model and data, but inherent surrogate errors need to be incorporated a priori. When we set $\sigma_{\text{sys}}=0$, i.e., turn off the systematic uncertainty entirely, we observe that the posteriors using the surrogate diverge from the posteriors using the direct afterglowpy evaluations. Although the surrogate posteriors still find values that are close the injection parameters, these parameters are not contained in the 95% credible limits. Likewise, the credibility contours between the surrogate and direct afterglowpy posterior do not overlap anymore. Thus, adjusting for the surrogate uncertainty remains crucial.
In Appendix B we show similar injection-recoveries for the pyblastafterglow and possis surrogates. While for these models we can not compare the fiesta posterior to a posterior that evaluates the likelihood with the physical base model, we still find good recovery of the injected parameters.
However, a systematic assessment of whether injections are generally well recovered requires going beyond individual examples. For this reason, we run 200 injection-recoveries each for afterglowpy, pyblastafterglow, and possis. This way, we obtain the distribution of the injection values’ posterior quantiles across multiple inferences. The cumulative distribution of these quantiles can be visualized in a P-P plot. Figure 5 shows the resulting P-P plot for the GRB afterglow inferences, in Fig. 6 we provide the P-P plot for the possis injection recoveries. Overall, we find that the injection values are recovered well, as the injected value lies outside the posterior only in 1.3% of cases for the afterglowpy and pyblastafterglow injections and in 5.1% of cases for the possis injections. However, if the posteriors were unbiased, then, according to the probability integral transform, the quantiles would adhere to a uniform distribution. In Fig. 5 and 6, it is apparent that this is not always the case and the cumulative distributions sometimes fall outside the grey 68%, 95%, or 99.7% confidence ranges in which they would fall if they were uniform.
There are several reasons for this behavior. In certain cases, the offset between the injection data with the physical base model and the surrogate used to sample the posterior might introduce biases in the recovery. However, in Fig. 5 and 6 we also show P-P plots for cases where the injection data is generated with the surrogate instead of the base model. These are shown as dashed lines in Fig. 5 and 6 and display the same trends as the P-P plots with the injections from the base model. Hence, we conclude that the suboptimal recovery of certain parameters is primarily due to our inclusion of the systematic uncertainty and the way we generate the mock data.
For the GRB afterglow injection-recoveries, it is the wing angle $\alpha_{w}$, the interstellar density $\log_{10}(n_{\rm ism})$, and the electron power index $p$ that show the most notable deviation from uniformity. In the case of $\alpha_{w}$, for instance, we note a high degeneracy with the output data. The lightcurve does not change noticeably when $\alpha_{w}$ goes from 2.5 to 3.5, unless perhaps the alignment with the observer changes, i.e., $\theta_{w}$ suddenly becomes larger than $\iota$. This leads to broad posterior support for $\alpha_{w}$, therefore causing the more extreme quantiles for $\alpha_{w}$ to be overrepresented in the P-P plot. This also applies to $\log_{10}(n_{\rm ism})$, which mostly affects the early part of the lightcurve prior to the jet-break. When $\log_{10}(n_{\rm ism})$ is low, this late part of the lightcurve will also match a jet with somewhat higher interstellar density and the posterior will have significant support above the injected value, overrepresenting low quantiles. Our mock injection data is spaced log-uniformly between 0.01 days to 200 days and thus in most cases will contain a large data segment from the post-jet break which may cause a degeneracy in $\log_{10}(n_{\rm ism})$. On the other hand, for the electron power index $p$, the injection value’s posterior quantile is too often too close to 0.5, i.e., $p$ is overdetermined. This can be attributed to the fact that the value for $p$ strongly influences the post-jet break slope from the data. The large segment of post-jet break data enables the sampler to infer $p$ with good accuracy, but the addition of the systematic uncertainty artificially broadens the posterior, which causes the injected value to lie too often at the center of the posterior.
A similar effect can be seen for some of the KN parameters in Fig. 6. The parameters $\bar{v}_{\rm ej,dyn}$, $\bar{Y}_{\rm e,dyn}$, $\bar{v}_{\rm ej,wind}$ show the same over-determination, i.e., quantiles too often lie close to $0.5$, when injecting with the possis base model. However, the corresponding P-P plots for injections from the surrogate show similar or even stronger over-determination across all parameters. When the injection is directly with possis, the parameters $\iota$, $\log_{10}(m_{\rm ej,dyn})$, $\log_{10}(m_{\rm ej,wind})$ also have an overrepresentation of lower quantiles. However, this can be attributed to the fact that when we inject possis data, we do so from randomly selected test data lightcurves in our set of fixed possis simulations. These were run on a discrete set of parameters and therefore, as seen for instance in the case of the inclination, the injected value will be exactly 00 rad instead of some small value if the injection value was drawn from a continuous distribution. Since 00 rad is the prior bound in the inference, the 00th quantile is overrepresented in the distribution of posterior quantiles. This also partially explains the higher percentage of $5.1$% mentioned above for the cases in which the injection value lies completely outside the posterior.
4.2 Performance
Figure 7: Sampling run time of a fiesta inference as a function of the effective sample size (ESS) with different numbers of sampling parameters. The plot shows the runtime of a pyblastafterglow injection-recovery when different numbers of nuisance parameters for the time-dependent systematic uncertainty are added. The performance test was conducted on two different GPU types as indicated by the colors in the legend.
The computational cost of sampling the posterior $P(\@vec{\theta}|d)$ is mainly determined by the cost of the likelihood function. Using ML surrogates reduces the evaluation time of the likelihood function by several orders of magnitude. This is best illustrated by comparing the total runtime of the inferences shown in Fig. 4, where the posterior for an afterglowpy injection was obtained with and without the surrogate. The total sampling time with fiesta amounts to $96$ s on an NVIDIA H100 GPU, whereas sampling with the actual afterglowpy model in nmma takes $19,700$ s ($\approx 5.5$ h) on 24 ${}^{\text{\textregistered}}$ ${}^{\text{\textregistered}}$ Silver CPUs. Using the power consumption values for the GPUs and CPUs used (NVIDIA Corporation, 2023, 2024; Intel Corporation, 2022), this implies that inferences with fiesta consume around $124$, respectively $168$ less energy than the equivalent CPU-based run when using the NVIDIA H100, respectively NVIDIA RTX 6000 GPU. This difference is mainly due to the speed-up in the likelihood evaluation and not related to the sampling algorithm, since the run that uses the fiesta surrogate within the nmma algorithm takes again 203 seconds on the same 24 CPUs. It should be noted that the flowMC sampler consists of several stages. First, the likelihood function and the associated neural networks are just-in-time compiled with jax, which in our inferences takes around $60$ s. Then, a training loop takes place which concurrently runs MCMC sampling and training of the normalizing flow proposal, which takes another $30$ s. The samples produced during the training loop are considered burn-in samples and are discarded. Generating the final set of posterior samples then only takes about $5$ s. The exact length of these different segments depends on the number of datapoints, the hyperparameters of flowMC, and the number of photometric filters in the data. The more filters there are, the more often the FluxModel output needs to be converted to AB magnitudes, which involves the evaluation of an integral, thus decelerating the likelihood evaluation.

### v2

@@ 3.3 KN surrogates
Figure 3: Benchmarks of two surrogates for the KN possis model. We show the deviations of surrogate predictions against a test data set of size $n_{\text{test}}=2238$. Figure layout as in Fig. 1. The figure compares two different surrogates: one using the MLP architecture (blue) and the other a LightcurveModel, where an MLP is trained for each passband separately (green).
fiesta includes surrogates for the KN model from the possis code (Bulla:2019muo; Bulla:2022mwo). possis is a 3D Monte Carlo radiation transport code that assumes homologously expanding ejecta and determines the emitted flux and polarization at each timestep from the Monte Carlo propagation of photon packets. As the photon packets move through the matter cells in the ejecta profile, they can interact with the matter through bound-bound and electron-scattering transitions, where the prescriptions take into account temperature-, density-, and electron-fraction dependent opacities (Tanaka:2019iqp). The BNS ejecta are assumed to follow a certain geometry inspired by numerical relativity simulations (Kiuchi:2017pte; Radice:2018pdn; Kawaguchi:2019nju; Hotokezaka:2018dfi). In particular, the ejecta have two components, the dynamical ejecta with mass $m_{\rm ej,dyn}$ and the wind ejecta with mass $m_{\rm ej,wind}$. The velocity in the dynamical component ranges from 0.1cc to a cutoff value that is determined such that the mass-averaged velocity is $\bar{v}_{\rm ej,dyn}$. Likewise, the wind component has the mass-averaged velocity $\bar{v}_{\rm ej,wind}$ with a minimum velocity of at least 0.02cc. The dynamical electron fraction varies with the polar angle $\theta$ according to
$Y_{e,\text{dyn}}=a\cos(\theta)^{2}+b\ ,$ (11)
where $a=0.71b$ (Setzer2023) and b is scaled to achieve the desired mass-averaged electron fraction $\bar{Y}_{e,\text{dyn}}$. This setup was also used in Anand:2023jbz; Ahumada:2025ubp. The electron fraction for the wind ejecta, however, is assumed to be constant and thus we do not mark its symbol $Y_{e,\text{wind}}$ with a bar.
We trained three possis surrogates, two as a FluxModel with an MLP and cVAE architecture, respectively, and the other as a LightcurveModel. These surrogates constitute an update for the Bu2019 surrogate implemented in nmma that was based on an older possis version. Our training data consists of $n_{\text{train}}=17899$ flux densities calculated from possis. For each of these computations we set the number of Monte Carlo photon packets to $10^{7}$. Parameters for the training data set are randomly drawn within the ranges from Table 3, though we note that the ejecta masses were uniformly in linear space, not in log-space. Furthermore, the inclination, $\iota$, is confined to a fixed grid from the possis output that is spaced uniformly in $\cos(\iota)$ between 0 (edge-on) and 1 (face-on). We also mention that the LightcurveModel additionally receives the redshift $z$ sampled randomly between 0 and 0.5 as a training parameter (the luminosity distance is fixed at 10 pc), so that the surrogate automatically incorporates the K-correction to the passband magnitudes. We find no significant performance difference for various hyperparameter settings, so the final number of PCA coefficients for the MLP architecture in the FluxModel and LightcurveModel was set to 100, the cVAE was trained on a down-sampled flux array of dimensions $64\times 40$. Training took about 4 minutes for the MLP architectures and 27 minutes for the cVAE on an H100 GPU.
In Fig. 3, we benchmark the surrogate models for the possis KN model. In particular, we compare the MLP FluxModel against the MLP LightcurveModel, since the FluxModel with the cVAE architecture underperforms both. This is because, in contrast to the GRB afterglow training data, the training data from possis contains inherent Monte Carlo noise, which is more difficult for the cVAE to “average out,” as it is trained directly on the flux output. The MLP architecture uses the coefficients of the principal components as training input and, thus, it is less sensitive to small noise fluctuations. Still, even for those architectures, we generally find higher deviations in the predictions from the test data set than for the afterglow surrogates. In particular, the mismatch rises drastically above 0.5 mag for many test data samples when the flux brightness suddenly drops. This can be seen in Fig. 3, where the mismatch in the right panels is typically confined to $\lesssim 0.5$ mag, but then spikes around the time the KN starts to fade (i.e., after around 1 day) in the UV band and after around $\gtrsim$ 10 days in the $i$-band. In general, the dimmer the light curve, the higher the Monte Carlo noise contribution and the larger the mismatch between the surrogate and the test data becomes. However, inspection of random test samples reveals that the surrogate prediction matches quite well at early times, when the absolute magnitude is still brighter than 10-10 mag. At very late times, we truncated the training data below $10^{6.5}$ mJys (corresponding roughly to the 0th absolute magnitude), which causes the mismatch to decrease again when the surrogates pick up on this trend. The LightcurveModel performs slightly better at these later times, however, the FluxModel performs better at earlier times when the emission is still bright. We thus set the MLP FluxModel as the default KN surrogate for the subsequent analyses in Sects. 4 and 5, as the early and bright parts of the light curve are most relevant for real-life applications.
We also note that the surrogates from Table 3 are rather extensive in their scope in which they can predict the spectral flux density. Specifically, the GRB afterglow surrogates have been trained across $10^{9}$$5\cdot 10^{19}$ Hz, i.e., from the radio to hard X-ray, and over a time interval of $10^{-4}$$2000$ days. The KN surrogates can predict the expected emission between 0.20.22626 days in the frequency range of $10^{14}$$2\cdot 10^{15}$ Hz, i.e., from the far infrared to the far ultraviolet (UV). In certain use cases, the surrogates could easily be retrained on a smaller frequency or time interval, to potentially deliver even better performance.
4 Bayesian inference with fiesta
Using the surrogate for the determination of the expected light curve $m^{\star}(t,\@vec{\theta})$ in Eq. (3) is an approximation to the real likelihood function. In this section, we demonstrate that this approximation is still capable of recovering the correct posterior when accounting for the surrogate uncertainty. We do so by using the best-performing surrogates presented in Sect. 3, namely, the FluxModel instances with cVAE architecture for the afterglow models, and the FluxModel with the MLP architecture for the KN model. We also discuss the performance of the flowMC implementation in fiesta and how it scales when the dimension of the parameter space is increased to include more systematic nuisance parameters in the sampling.
4.1 Injection recoveries
Figure 4: Parameter recovery for an injected mock light curve from the gaussian afterglowpy jet model. The corner plot shows the posterior contours at 68% and 95% credibility. Parameters correspond to the symbols in Table 3, $\sigma_{\text{sys}}$ is the freely sampled systematic uncertainty. Different colors compare posteriors obtained with different sampling methods. The posterior in red is based on likelihood evaluations from the proper afterglowpy model with the nmma sampler. The purple posterior relies on the fiesta surrogate for the likelihood evaluation but uses the nmma sampler. The light blue posterior uses the fiesta surrogate as well but is sampled within fiesta’s own inference framework that relies on flowMC. The injection parameters used to generate the mock light curve data are indicated by the orange lines. The insets on the upper right side show the injection data across the photometric filters and the best-fit light curve (i.e., highest likelihood) of the fiesta posterior (lightblue) and the actual afterglowpy light curve used to generate the mock data (red). The latter lies almost completely underneath the former.
Figure 5: P-P plots for GRB afterglow injections. Each panel shows a P-P plot for the recovery of the parameter displayed in its top left corner. The P-P plots show the cumulative distribution of the injected values’ posterior quantiles for 200 injections. The lightblue curves indicate injection recoveries with afterglowpy, the magenta ones for pyblastafterglow. The solid lines signify that the injections stem from physical base model, the dashed lines indicate an injection with the surrogate itself. The gray areas mark the 68%-95%-99.7% confidence range in which the quantile distribution should fall if it was uniformly distributed.
To evaluate whether fiesta’s surrogates are capable of recovering the correct posterior, we created mock light curve data with the physical base model using randomly drawn model parameters. We were then able to obtain the posterior using the surrogate. The injection data always contain 75 mock magnitude measurements across multiple bands, encompassing the frequency and time range of the model and representing a well-sampled light curve. Specifically, for GRB afterglows, the injection data span from 0.01 to 200 days and contains mock observations from the radio to the X-ray bands. The KN injections reach from the infrared to UV and contain data points between 0.5 to 20 days. For the KN injections, we also apply a detection limit of 24 apparent mag at 40 Mpc, which prevents the surrogates of being used in regions with high prediction error due to high Monte Carlo noise in the possis training data. We add Gaussian noise to these mock measurements, where the measurement errors $\sigma(t_{j})$ are drawn from a $\chi^{2}$-distribution with one degree of freedom and scaled to lie around $0.1$ mag. To recover the posterior with the surrogate, we use uniform priors across the parameter ranges specified in Table 3. The luminosity distance is fixed to $40$ Mpc for each injection.
In Fig. 4, we show how the parameters of one particular injection from the afterglowpy Gaussian jet model are recovered with fiesta’s surrogates, using either PyMultinest in nmma or flowMC as sampler. Since the injection also incorporates a random mock measurement error, the posterior is not always centered around the true injected parameters indicated by the orange lines, but the marginalized posteriors contain the injected values in the 95% credible interval.
Of all the models presented in Sects. 3.2 and 3.3, afterglowpy is the only one where the execution time of a single likelihood call is sufficiently fast to be used directly in Bayesian inference. Thus, we can compare the approximate posterior obtained with the fiesta surrogate to the posterior obtained using the actual physical base model for the likelihood evaluation. The latter is shown in red in Fig. 4. We find good agreement with the posteriors obtained with the fiesta surrogate, though some small deviations in the posterior tails for the inclination and jet opening angle exist.
We also ran four additional afterglowpy injection recoveries (similar to those in Fig. 4) and compared them to the surrogate posteriors, finding good agreement. For certain degenerate parameters, we observed that even the flowMC sampling algorithm with the surrogate recovers the true injected values better; namely, the true value lies more at the center of the posterior than with nmma’s PyMultinest sampler and when using the actual afterglowpy model for the likelihood. This can be attributed to broader exploration in the parameter space for degenerate parameters such as $\epsilon_{B}$ and $\log_{10}(n_{\rm ism})$, indicating an advantage in flowMC’s dependence on gradient-based samplers and global proposals.
To systematically evaluate the disagreement between the posteriors caused by the surrogates, we resorted to the Kullback-Leibler (KL) divergence $D_{KL}$. Bevins:2025srk assessed a theoretical link between the root mean square error (RMSE) of an emulator and the impact on the posterior in terms of $D_{KL}$ if it was obtained with the base model or with the surrogate. For a linear model, they derive the upper-bound on $D_{KL}$,
$D_{KL}\leq\frac{N_{d}}{2}\frac{\rm{RMSE}^{2}}{\sigma^{2}}\ ,$ (12)
where $N_{d}$ is the number of data points and $\sigma$ their typical error scale. While this upper bound assumes a linear relation $d(\@vec{\theta})$, we can still examine this relationship using our afterglowpy injections and the respective surrogate and base model posteriors. We then determine $D_{KL}$ between the posteriors through a kernel density estimate. Using our 5 injection-recoveries with $N_{d}=75$, we find that $D_{KL}$ ranges between 2 – 8 nats and the upper limit is indeed obeyed when setting $\rm{RMSE}=0.1$ mag and taking the largest error from the injected mock data as a conservative value for $\sigma$. In fact, Eq. (12) is a factor of 1.4 – 2.8 above our estimated value for $D_{KL}$. It should be noted however, that this only serves as a limited sanity check, since the bound in Eq. (12) is derived under simplifying assumptions, and determining $D_{KL}$ through kernel density estimates might not be numerically accurate.
We also point out that in order to achieve this agreement between the surrogate posterior and the posterior based on afterglowpy directly, we set a minimum threshold on the systematic uncertainty $\sigma_{\text{sys}}$. Specifically, $\sigma_{\text{sys}}$ was sampled as a free parameter with a uniform prior $\sigma_{\text{sys}}\sim\mathcal{U}(0.3,1)$. Lowering the limit on $\sigma_{\text{sys}}$ can lead to biases in the posteriors recovered with the surrogate, since sampling the systematic uncertainty mainly accounts for potential tension between model and data, but inherent surrogate errors need to be incorporated a priori. When we set $\sigma_{\text{sys}}=0$, i.e., turn off the systematic uncertainty entirely, we observe that the posteriors using the surrogate diverge from the posteriors using the direct afterglowpy evaluations. Although the surrogate posteriors still find values that are close the injection parameters, the latter are not contained in the 95% credible limits. Likewise, the credibility contours between the surrogate and direct afterglowpy posterior do not overlap anymore. Thus, adjusting for the surrogate uncertainty remains crucial.
In Appendix B, we show similar injection recoveries for the pyblastafterglow and possis surrogates. For these models, we could not compare the fiesta posterior to a posterior that evaluates the likelihood with the physical base model, yet we still found a good recovery of the injected parameters.
However, a systematic assessment of whether injections are generally well recovered requires going beyond individual examples. For this reason, we ran 200 injection recoveries each for afterglowpy, pyblastafterglow, and possis. This way, we obtained the distribution of the injection values’ posterior quantiles across multiple inferences. The cumulative distribution of these quantiles can be visualized in a P-P plot. Figure 5 shows the resulting P-P plot for the GRB afterglow inferences, in Fig. 6 we provide the P-P plot for the possis injection recoveries. Overall, we find that the injection values are recovered well, with the injected values lying within the posterior samples in 98.7% of the afterglowpy and pyblastafterglow injections and in 94.8% the possis injections. However, if the posteriors were unbiased, then, according to the probability integral transform, the quantiles would adhere to a uniform distribution. In Figs. 5 and 6, it is apparent that this is not always the case and the cumulative distributions sometimes fall outside the grey 68%, 95%, or 99.7% confidence ranges in which they would lie if they were uniform.
There are several reasons for this behavior. In certain cases, the surrogate error might introduce biases in the recovery. However, in Figs. 5 and 6 we also show P-P plots for cases where the injection data is generated with the surrogate instead of the base model. These are shown as dashed lines in Figs. 5 and 6 and display the same trends as the P-P plots with the injections from the base model. Hence, we conclude that the suboptimal recovery of certain parameters is primarily due to our inclusion of the systematic uncertainty and the way we generate the mock data.
For the GRB afterglow injection recoveries, it is the wing angle $\alpha_{w}$, the interstellar density $\log_{10}(n_{\rm ism})$, and the electron power index $p$ that show the most notable deviation from uniformity. In the case of $\alpha_{w}$, for instance, we note a high degeneracy with the output data. The light curve does not change noticeably when $\alpha_{w}$ goes from 2.5 to 3.5, unless perhaps the alignment with the observer changes, i.e., $\theta_{w}$ suddenly becomes larger than $\iota$. This leads to broad posterior support for $\alpha_{w}$, therefore causing the more extreme quantiles for $\alpha_{w}$ to be overrepresented in the P-P plot. This also applies to $\log_{10}(n_{\rm ism})$, which mostly affects the early part of the light curve prior to the jet-break. When $\log_{10}(n_{\rm ism})$ is low, this late part of the light curve will also match a jet with somewhat higher interstellar density and the posterior will have significant support above the injected value, overrepresenting low quantiles. Our mock injection data are spaced log-uniformly from 0.01 days to 200 days and thus in most cases will contain a large data segment from the post-jet break which may cause a degeneracy in $\log_{10}(n_{\rm ism})$. On the other hand, for the electron power index $p$, the injection value’s posterior quantile is too often too close to 0.5, i.e., $p$ is overdetermined. This can be attributed to the fact that the value for $p$ strongly influences the post-jet break slope from the data. The large segment of post-jet break data enables the sampler to infer $p$ with good accuracy, but the addition of the systematic uncertainty artificially broadens the posterior, which causes the injected value to lie too often at the center of the posterior.
A similar effect can be seen for some of the KN parameters in Fig. 6. The parameters $\bar{v}_{\rm ej,dyn}$, $\bar{Y}_{\rm e,dyn}$, and $\bar{v}_{\rm ej,wind}$ show the same over-determination, i.e., their quantiles lie too often close to $0.5$. However, the corresponding P-P plots where injection data is constructed with the surrogate show similar or even stronger over-determination across all parameters. When the injection is directly with possis, the parameters $\iota$, $\log_{10}(m_{\rm ej,dyn})$, $\log_{10}(m_{\rm ej,wind})$ also have an overrepresentation of lower quantiles. However, this can be attributed to the fact that when we inject possis data, we do so from randomly selected test data light curves in our set of fixed possis simulations. These were run on a discrete set of parameters and therefore, as seen for instance in the case of the inclination, the injected value will be exactly 0 rad instead of some small value if the injection value was drawn from a continuous distribution. Since 0 rad is the prior bound in the inference, the 0th quantile is overrepresented in the distribution of posterior quantiles. This also offers a partial explanation for the 5.1% of injections mentioned above, where the posterior samples lie exclusively above or below the injection value.
4.2 Performance
Figure 7: Sampling run time of a fiesta inference as a function of the parameter space dimension. The plot shows the runtime of a pyblastafterglow injection recovery per effective sample size (ESS) when different numbers of nuisance parameters for the time-dependent systematic uncertainty are added. The performance test was conducted on two different GPU types as indicated by the colors in the legend.